The shaft of a motor rotates at a constant angular speed of 360rev/min. Angle
turned through in 1 sec in radian is? A. ^
B. 3^ C. 6^
D. 12^
(^=pie)
1
Expert's answer
2021-09-24T09:26:33-0400

360 revolutions per minute is 6 revolutions per second (360/60=6). Six revolutions are 6 times the angle of a circle (360° or 2·pi), therefore, the angle is


"\phi=6·2\pi=12\pi."